I need k answered step by step.

k. What is the value of a perpetual bond with an annual coupon of $100 if is required rate of return is 10 percent? 13 percent? 7 percent? Assess the fol- lowing statement: Because perpetual bonds match an infinite investment horizon, they have little interest rate price risk.
Answer #1

Value of perpetual bond = Annual coupon/ Rate of interest

1: Value of perpetual bond= 100/10% = $1000

2: Value of perpetual bond = 100/13% = $769.23

3: Value of perpetual bond = 100/7% = 1428.57

The statement is false because the perpetual bond is subject to high interest rate risk due to its high maturity period.